The length of a rectangle is 4cm, longer than its width.
If the perimeter of the rectangle is 32cm , find its area.

Answer by edjones(7569)

L=w+4
2(L+w)=p
2(w+4+w)=32 Substitute w+4 for L.
2(2w+4)=32
4w+8=32
4w=24
w=6
L=10
.
Lw=A
10*6=60 cm^2
.
Ed